OpenStudy (anonymous):

How do i go about this rational expression problem? Bob and Jack enter a 1000-km bicycle race. Bob cycles 4 km/h faster than Jack, but his bicycle gets a flat tire, which takes 30 minutes to repair. If the both of them finish the race in a tie, then how fast was each boy cycling? Give your answer in one decimal place.

OpenStudy (anonymous):

assume the velocity for jack=v then velocity for bob=v+4 total time for bob=\[[1000/(4+v)]+0.5\] it has to equal to time for jack=\[1000/v\] after solving the 2 equition, we got v=87.465, and v+4=91.465 for only one decimal v=jack=87.5 km/h v+4=bob=91.5 km/h

OpenStudy (anonymous):

How would i go about actually solving that equation though?

OpenStudy (anonymous):

\[[1000/(4+v)]+0.5=1000/v\] \[(1000+2+0.5v)/(4+v)=1000/v\] times the button to the other side \[1002v+0.5v^2=4000+1000v\]\[0.5v^2+2v-4000=0\] use the formula v= \[[-b+\sqrt{b^2-4ac}]/2a\]